Solve the equation for x
ln(3x+1)=ln(2x-8)

Answer by Alan3354(30959)

Solve the equation for x
ln(3x+1)=ln(2x-8)
------------
It the logs are equal, then
3x+1 = 2x-8
x = -9
----------
But, this give the log(-26) which does not exist.
--> no solution